a) Degree of operating leverage = Contribution margin / profit

Degree of operating leverage = 15,600,000 / 6,000,000 = 1.625

Sales = 1200000 x 24 = 28800000

Less: Variable cost = 1200000 x 11 = 13200000

Contribution Margin = 28800000 - 13200000 = 15,600,000

Less: Fixed cost = 9,600,000

Profit = 15600000-9600000 = 6000000

b) Break even units

Contribution margin per unit = Selling price - variable cost = $24 - $11 = $13

Break even units = Fixed cost / contribution margin per unit

Break even units = 9,600,000 / $13 = 738,461.5 or 738,462 units

c) If sales units increase by 25%

The new sales will be 1200,000 x 1.25 = 1500,000 cans

The new operating profit = [1500,000 x ($24-$11)] - $9,600,000 = $9,900,000

Increase in operating income in Dollars = $9,900,000 - $6,000,000 = $3,900,000

Percentage change in profits = 3,900,000 / 6,000,000 = 65% increased

Degree of operating leverage = Change in profit / change in sales

Degree of operating leverage = 65% / 25% = 2.6

D) If the company spends $30,000 as additional advertising expense (fixed cost) sales volume will increase by 10%. determine the new degree of operating leverage and new breakeven point in units
Answer #1

Correct Answer:

Breakeven point in units

740,769 units

Degree of operating leverage

2.28 Times

Working:

Units sold = 12000000 + (10% * 12000000)

= 1320000 units

Per unit

1320000 UNITS

A

Sales

$                  24

$         3,16,80,000

B

Less: Variable cost

$                  11

$         1,45,20,000

C =A-B

Contribution margin

$                  13

$         1,71,60,000

D

Less: Fixed cost

$   9,630,000

$            9,630,000

E =C-D

Net income

$            75,30,000

F =D/C

Breakeven point in units

740769.2

G =C/E

Degree of operating leverage

2.28

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.